if n people attened a social gathering and each person shakes hands with each of the other peopleexactly once, then the number of handshakesthat will take place is give by the formula n(n-1)= 36 handskes occured?
/2

Answer by venugopalramana(3286)

if n people attened a social gathering and each person shakes hands with each of the other peopleexactly once, then the number of handshakesthat will take place is give by the formula n(n-1)= 36 handskes occured?
FORMULA IS NC2
THAT IS N(N-1)/2=36
N(N-1)=72
HENCE N=9 BY INSPECTION SINCE 9*(9-1)=9*8=72